# Entanglement generation and scaling from noisy quenches across a quantum critical point

### v2

@@ Abstract
We study the impact of noise on the dynamics of entanglement in the transverse-field Ising chain, with the field quenched linearly across one or both of the quantum critical points of the model. Taking concurrence as a measure of entanglement, we find that a quench generates entanglement between nearest- and next-nearest-neighbor spins, with noise reducing the amount of entanglement. Focussing on the next-nearest-neighbor concurrence, known to exhibit Kibble-Zurek scaling with the square root of the quench rate in the noiseless case, we find a different result when noise is present: The concurrence now scales logarithmically with the quench rate, with a noise-dependent amplitude. This is also different from the “anti-Kibble-Zurek” scaling of defect density with quench rate when noise is present, suggesting that noisy entanglement generation is largely independent from the rate of defect formation. Intriguingly, the critical time scale beyond which no entanglement is produced by a noisy quench scales as a power law with the strength of noise, with the same exponent as that which governs the optimal quench time for which defect formation is at a minimum in a standard quantum annealing scheme.
@@ I Introduction
The study of quantum phase transitions (QPTs) form a core area of many-body physics. Such transitions - between ground states of different symmetry [1] or topology [2] - play out at zero temperature by varying a control parameter. As markers, physical observables exhibit scaling and universality when approaching quantum criticality, reflecting the divergence of intrinsic length- and time scales.
@@
Central to KZM is a universal scaling law that governs how the density of defects $n_{0}$ scales with the rate (inverse time scale) $1/\tau$ of a ramped quench (a.k.a. linear quench) across a continuous QPT: $n_{0}\sim\tau^{-\beta}$ with $\beta=d\nu/(1+z\nu)$, where $\nu$ and $z$ are the critical correlation and dynamic exponents respectively, and $d$ is the dimensionality of the system. The subindex on $n_{0}$ signifies that the quench is noiseless.
@@
The inverse of the defect density determines the average length $\lambda$ of the ordered domains in the broken-symmetry phase. The same length $\lambda$ appears to control also the amount of entanglement generated by the quench, suggesting a connection between defect formation and entanglement production. Evidence for this in the case of a ramped quench - from paramagnetic to ferromagnetic phase of the transverse-field Ising chain - was given by Cincio et al. in Ref. [19], confirming and extending results in [14]. Cincio et al. showed that the amount of entanglement entropy of a block with $L$ spins generated by a quench saturates at $\approx\frac{1}{6}\log_{2}\sqrt{\tau}$ when $L$ is large. Knowing that the block entanglement in one dimension saturates at $\frac{1}{6}\log_{2}\lambda$ with $\lambda<L$ an intrinsic length scale [32], this is in accord with the KZM scaling law that predicts that $\lambda$ for the quenched transverse-field Ising chain is given by $\lambda\sim\sqrt{\tau}$. The authors propose that the scenario is generic, and that the process that determines the defect density in the out-of-equilibrium dynamics of a QPT is also responsible for the entanglement that is generated.
@@
A different take was presented by Sengupta and Sen in Ref. [33]. These authors probed the two-spin entanglement generated by a power-law quench across both quantum critical points $h=h_{c}=\pm 1$ in the XY chain with a transverse field $h$, with the quench starting [ending] deep in a paramagnetic phase, with $h_{\text{initial}}\rightarrow$-$\infty$ $h_{\text{final}}\rightarrow\infty]$. It was found that when the time scale $\tau$ of the quench is larger than a critical value, $\tau>\tau_{0}$, only spins separated by an even number of lattice spacings get entangled, while for $\tau<\tau_{0}$ there is no two-spin entanglement; instead the entanglement becomes multipartite. In the first case the concurrence that measures the two-spin entanglement scales as $\sqrt{\alpha/\tau}$, where $\alpha$ is the exponent of the power-law quench. Again, the entanglement production for a slow quench is seen to scale with the square root of the quench rate, similar to the defect density predicted by KZM for the Ising and XY chains with a transverse field. Additional results on the transverse-field XY chain, pertaining also to entanglement negativity and quantum discord are in line with this result [34, 35], with related findings in Ref. [36]. Subsequent works have explored how the decoherence processes that compete with the entanglement generation scale with the quench rate [37, 38].
@@
The analytical and numerical findings in Refs. [14, 19, 33, 34, 35], although limited to simple one-dimensional models, indeed suggest that a quench across a QPT, no matter how slow, will invariably lead to entanglement generation. Moreover, the KZM-type scaling of entanglement, as measured by concurrence, negativity and quantum discord indicates that the break-up of an ordered phase by defects lies at the very heart of the process. A simple argument that directly links the scaling of concurrence to that of defect formation is contained in the analysis in Ref. [34]. Given the nonlocality of entanglement, one may have thought that the scaling of entanglement generation with $\tau$ should rather be governed by density correlators which, as shown in Refs. [39, 40], cannot be derived solely from the assumptions underlying KZM, and hence go beyond the scaling of defect density formation. Importantly, however, and as shown explicitly in Ref. [39] for the two-point density correlators, the relevant time- and length scales that appear in these correlators do obey KZM scaling. This resolves a potentially thorny conceptual issue.
@@
The first steps in this direction were taken by Dutta et al. [50] who showed that an isolated system driven across a QPT by a noisy control field exhibits an “anti-Kibble-Zurek” behavior, whereby slower driving results in a higher density of defects. The authors argued that the KZM dynamics decouples from noise-induced effects, leading to an additive scaling formula for the density of defects in the limit of weak noise, $n\sim r\tau+c\tau^{-\beta}$, with $r$ the rate at which the presence of noise generates defects, and with $c$ a nonuniversal constant. When the first term dominates, the anti-Kibble-Zurek behavior becomes manifest, as fleshed out in later theoretical [51, 52, 53, 54, 55, 56] and experimental [57] studies (cf. Appendix A).
@@
We here address these questions in the setting of the transverse-field Ising (TFI) chain, the same model as used in the case studies in Refs. [50, 51]. Specifically, we analyze the two-spin entanglement generated by noisy quenches across the QPTs of the model, taking white noise as a representative, and added as a classical stochastic variable to the transverse field. Since one does not have access to individual noise realizations in an experiment, we consider the ensemble average over many realizations, with the average density matrix $\langle\rho\rangle$ emerging as the relevant quantity of interest. To obtain $\langle\rho\rangle$, we solve an exact master equation for the quench dynamics averaged over the noise distribution. By numerically solving the master equation we then extract the average entanglement produced by an ensemble of noisy quenches. When two-spin entanglement is present, as measured by concurrence, we find that it scales logarithmically with the time scale $\tau$ when the quenches are slow, with an amplitude that depends on the intensity $\xi$ of the corresponding white-noise distribution. This is different from the behavior of the defect density in the presence of white noise as predicted by the anti-Kibble-Zurek additive formula in Refs. [50, 51]. As a consequence, the notion that entanglement and defect formation are intrinsically linked becomes less persuasive. A possible counterpoint is our finding that the time scale $\tau_{c}(\xi)$ beyond which no two-spin entanglement is produced scales as a power law of the noise intensity with the very same exponent as that of the optimal quench time for which the density of defects is minimized in a quantum annealing scheme [50, 51]. Whereas a direct, quantitative link between two-spin entanglement generation and defect formation seems unlikely in the presence of noise, this latter result hints at a possible connection, albeit more complex, between the two phenomena.
To prepare for our analysis, in the next section we set up the formalism and define the model. As a backdrop we then revisit the simple case of a noiseless quench in Sec. III. This allows us to present a new result: Different from a noiseless symmetric “very-large-field” quench as discussed in Ref. [33] (i.e., with $h_{i}=-h_{f}\rightarrow$-$\infty$), we find that a generic quench generates two-spin entanglement also between nearest-neighbor spins. In Sec. IV we turn to the main focus of our work, the case of noisy quenches. We take off from the exact master equation mentioned above, display the results from its numerical solution, and extract the scaling laws for the two-spin averaged entanglement as measured by concurrence - generated by slow noisy quenches across the quantum critical points of the model. Sec. V contains a brief summary and outlook. Mathematical details of our analysis can be found in the Appendix.
@@ II Model and formalism
(where $s^{\pm}_{n}=\sigma^{\pm}_{n}/2=(\sigma^{x}_{n}\pm i\sigma^{y}_{n})/2$), and introducing the discrete Fourier transforms $c_{k}=(1/\sqrt{N})\sum_{n}e^{ikn}c_{n}$ with $k=\pm(2j\!-\!1)\pi/N,j=1,2,\ldots N/2$, $H_{0}(t)$ can be expressed as a sum over decoupled mode Hamiltonians $H_{0,k}(t)$,
@@
set by the angle $\theta_{k}(t)=\arctan({\Delta_{k}}/{h_{0,k}(t)})/2$. In this basis the Hamiltonian in Eq. (3) takes the diagonal form
@@
with $\varepsilon_{k}(t)=\sqrt{h^{2}_{k}(t)+\Delta_{k}^{2}}$. The eigenstates of $H_{0}(t)$ factorize into quasiparticle states labelled by $k$. Introducing $|0\rangle$ as the vacuum of the JW fermions $\{c_{k},c_{k}^{\dagger}\}$, these states can be written as
@@
In the case of a noiseless ramped quench we can write $h_{0}(t)\!=\!(t-t_{i})/\tau+h_{i}$, with $\tau=(t_{f}-t_{i})/(h_{f}-h_{i})$ the time scale of the quench, where $h_{i}\,(h_{f})$ is the initial (final) value of the ramp field at time $t_{i}\,(t_{f})$. To address the problem of noisy quenches, a direct approach would be to add a stochastic variable $\eta(t)$ to the ramp field, making the replacement $h_{0}(t)\rightarrow h(t)=h_{0}(t)+\eta(t)$ in the formulas above, compute the observable of interest - or, in our case, an entanglement measure - and then take an average over many noise realizations $\eta(t)$. In practice this is a cumbersome procedure, which, however, can be cut short by using an exact master equation, to be discussed in Sec. IV (see also Appendix C.)
A standard choice for this class of applications is to take the noise to be Gaussian with vanishing mean, $\langle\eta({t})\rangle=0$, and with Ornstein-Uhlenbeck two-point correlations [61]
@@
Here $\tau_{n}$ is the noise correlation time and $\xi$ the noise amplitude for fixed $\tau_{n}$. For simplicity, and for comparison with the results in Refs. [50, 51, 52, 53, 54, 55, 62], we shall confine our attention to the common white-noise limit,
@@ III Entanglement generation from a noiseless ramped quench
In Fig. 1(a), we present our results for the concurrence $C_{\l ,\l +1}$ between two nearest-neighbor spins as a function of the endpoint quench field $h_{f}=h_{0}(t)$ at time $t$, with the quench starting at time $t_{i}$ with $h_{i}=-30$. Three graphs are plotted, for a slow ($\tau=500$), intermediate ($\tau=10$), and fast ($\tau=0.2$) quench, all for a chain with $N=200$ sites. The black dotted curve serves as a reference, showing $C_{\l ,\l +1}$ for the instantaneous ground state. As anticipated, and as shown in the inset of Fig. 1(a), this curve is in agreement with that known in the literature for $h_{0}(t)\geq 1/2$ [64, 65]. Its minimum at $h_{0}(t)=0$ reflects the multipartite entanglement of the zero-field GHZ ground state for a finite ferromagnetic TFI chain, implying a suppression of the two-spin entanglement.
@@
For intermediate values of the time scale, with $\tau$ in the interval $1\lesssim\tau\lesssim 100$, $C_{\l ,\l +1}$ exhibits oscillations, exemplified by the green curve in the inset of Fig. 1(a). The oscillations, originating from the phase factors of the off-diagonal elements ($v_{0,k}(t)u_{0,k}^{\ast}(t)$, $v_{0,k}^{\ast}(t)u_{0,k}(t)$) of the density matrix $\rho_{0,k}(t)$ [14, 39], become more volatile in the lower range of the interval $1\lesssim\tau\lesssim 100$, but get dampened as $h_{0}(t)$ increases.
As seen from the blue graph for $\tau=0.2$, a further reduction of the time scale of the quench causes $C_{\l ,\l +1}$ to abruptly drop to zero after crossing the first quantum critical point at $h_{0}(t)=-1$. $C_{\l ,\l +1}$ then suddenly shoots up when crossing the second critical point at $h_{0}(t)=1$, with one more spike soon thereafter before vanishing identically. This striking behavior, common to all time scales $\tau\lesssim 0.5$, shows that a very fast quench across a QPT may completely wipe out the two-spin entanglement, instead triggering multipartite entanglement (cf. the instantaneous GHZ ground state at $h_{0}(t)=0$) or a mutually factorized two-spin state, with the latter possibly forming an entangled composite with other spins of the system.
@@
Figure 2: (Color online) Noise-averaged concurrence between two spins at time $t$ for a ramped quench of time scale $\tau=10$ with white noise of different intensities $\xi$ added to the transverse field of the TFI chain, with quench interval from $h_{i}=-30$ at $t=t_{i}$ to $h_{f}=h_{0}(t)$: (a) nearest-neighbor spins, and (b) next-nearest-neighbor spins. System size: $N=200$.
@@
Figure 3: (Color online) Noise-averaged concurrence between two spins as a function of time scale $\tau$ for a ramped quench with added white noise of different intensities $\xi$, with quench interval from $h_{i}=-30$ to (a) $h_{f}=5$ for nearest-neighbor spins; (b) $h_{f}=30$ for next-nearest-neighbor spins. System size: $N=200$.
Independent of the choice of time scale $\tau$, all quenches generate entanglement between neighboring spins which decays to zero when $h_{0}(t)\rightarrow\infty$, consistent with the analytical prediction by Sengupta and Sen [33]. These authors predicted a different type of behavior for the next-nearest-neighbor concurrence $C_{\l ,\l +2}$, with two distinct asymptotics for very large $h_{0}(t)$: $C_{\l ,\l +2}\neq 0$ $C_{\l ,\l +2}=0$] when $\tau>\tau_{0}\approx 1.96$ $\tau<\tau_{0}$]. Our results for $C_{\l ,\l +2}$, displayed in Fig. 1(b) as a function of the endpoint quench field $h_{0}(t)$ at time $t$, with the quench starting at time $t_{i}$ with $h_{i}=-30$, agrees with this prediction: By extrapolating to large $h_{0}(t)$, we find numerically that $1.9<\tau_{0}<2.0$.
While consistent with the prediction of the time scale $\tau_{0}$ [33], the overall behavior of $C_{\l ,\l +2}$ as shown in Fig. 1(b) is unexpected from the perspective of KZM. Only for slow quenches (cf. the red curve with $\tau=500$ in Fig 1(b)) is there a weak imprint on $C_{\l ,\l +2}$ that the quench has approached or crossed the critical point $h_{c}=-1$ (but not $h_{c}=1$). Besides this, values of $C_{\l ,\l +2}$ remain very small - not visible in Fig. 1(b) - for all $h_{0}(t)\lesssim 2$. This contrasts with the nearest-neighbor concurrence $C_{\l ,\l +1}$ which exhibits a dramatic increase when approaching the first critical point $h_{c}=-1$, supporting the conjecture that the entanglement generated reflects the KZM formation of defects when the dynamics turns nonadiabatic [19, 14]. One may naively have thought that the fast drop of $C_{\l ,\l +1}$ in the ferromagnetic phase $1<h_{0}(t)<1$, cf. Fig. 1(a), would influence $C_{\l ,\l +2}$ by a restructuring of the entanglement with closest further neighbor spins. But not so. Instead, the next-nearest-neighbor concurrence is driven by the postcritical dynamics when the magnitude of the transverse field becomes sufficiently large. It is interesting to note that for near-adiabatic evolution (very slow quench: $\tau=500$), $C_{\l ,\l +2}$ closely follows its instantaneous counterpart (static reference curve) only when $h_{0}(t)<h_{c}=-1$, as illustrated in Fig. 1(b). Despite expectations, its behavior is distinct from the reference curve for $h_{0}(t)>-1$. This finding, similar to that above for $C_{\l ,\l +1}$ with $h_{0}(t)\gtrsim 1$, reflects how entanglement generation in non-equilibrium systems can differ from its equilibrium counterpart, even during a nearly adiabatic evolution.

### v3

@@ Abstract
We study the impact of noise on the dynamics of entanglement in the transverse-field Ising chain, with the field quenched linearly across one or both of the quantum critical points of the model. Taking concurrence as a measure of entanglement, we find that a quench generates entanglement between nearest- and next-nearest-neighbor spins, with noise reducing the amount of entanglement. Focusing on the next-nearest-neighbor concurrence, known to exhibit Kibble-Zurek scaling with the square root of the quench rate in the noiseless case, we find a different result when noise is present: The concurrence now scales logarithmically with the quench rate, with a noise-dependent amplitude. This is also different from the “anti-Kibble-Zurek” scaling of defect density with quench rate when noise is present, suggesting that noisy entanglement generation is largely independent of the rate of defect formation. Intriguingly, the critical time scale beyond which no entanglement is produced by a noisy quench scales as a power law with the strength of noise, with the same exponent as that which governs the optimal quench time for which defect formation is at a minimum in a standard quantum annealing scheme.
@@ I Introduction
The study of quantum phase transitions (QPTs) form a core area of many-body physics. Such transitions - between ground states of different symmetry [1] or topology [2] - play out at zero temperature by varying a control parameter. As markers, physical observables exhibit scaling and universality when approaching quantum criticality, reflecting the divergence of intrinsic length and time scales.
@@
Central to KZM is a universal scaling law that governs how the density of defects $n_{0}$ scales with the rate (inverse time scale) $1/\tau$ of a ramped quench (also known as linear quench) across a continuous QPT: $n_{0}\sim\tau^{-\beta}$ with $\beta=d\nu/(1+z\nu)$, where $\nu$ and $z$ are the critical correlation and dynamic exponents respectively, and $d$ is the dimensionality of the system. The subindex on $n_{0}$ signifies that the quench is noiseless.
@@
The inverse of the defect density determines the average length $\lambda$ of the ordered domains in the broken-symmetry phase. The same length $\lambda$ appears to control also the amount of entanglement generated by the quench, suggesting a connection between defect formation and entanglement production. Evidence for this in the case of a ramped quench - from paramagnetic to ferromagnetic phase of the transverse-field Ising (TFI) chain - was given by Cincio et al. in Ref. [19], confirming and extending results in Ref. [14]. Cincio et al. showed that the amount of entanglement entropy of a block with $L$ spins generated by a quench saturates at $\approx\frac{1}{6}\log_{2}\sqrt{\tau}$ when $L$ is large. Knowing that the block entanglement in one dimension saturates at $\frac{1}{6}\log_{2}\lambda$ with $\lambda<L$ an intrinsic length scale [32], this is in accord with the KZM scaling law that predicts that $\lambda$ for the quenched transverse-field Ising chain is given by $\lambda\sim\sqrt{\tau}$. The authors propose that the scenario is generic, and that the process that determines the defect density in the out-of-equilibrium dynamics of a QPT is also responsible for the entanglement that is generated.
@@
A different take was presented by Sengupta and Sen in Ref. [33]. These authors probed the two-spin entanglement generated by a power-law quench across both quantum critical points $h=h_{c}=\pm 1$ in the XY chain with a transverse field $h$, with the quench starting [ending] deep in a paramagnetic phase, with $h_{\text{initial}}\rightarrow$-$\infty$ $h_{\text{final}}\rightarrow\infty]$. It was found that when the time scale $\tau$ of the quench is larger than a critical value, $\tau>\tau_{0}$, only spins separated by an even number of lattice spacings get entangled, while for $\tau<\tau_{0}$ there is no two-spin entanglement; instead the entanglement becomes multipartite. In the first case, the concurrence that measures the two-spin entanglement scales as $\sqrt{\alpha/\tau}$, where $\alpha$ is the exponent of the power-law quench. Again, the entanglement production for a slow quench is seen to scale with the square root of the quench rate, similar to the defect density predicted by KZM for the Ising and XY chains with a transverse field. Additional results on the transverse-field XY chain, pertaining also to entanglement negativity and quantum discord are in line with this result [34, 35], with related findings in Ref. [36]. Subsequent works have explored how the decoherence processes that compete with the entanglement generation scale with the quench rate [37, 38].
@@
The analytical and numerical findings in Refs. [14, 19, 33, 34, 35], although limited to simple one-dimensional models, indeed suggest that a quench across a QPT, no matter how slow, will invariably lead to entanglement generation. Moreover, the KZM-type scaling of entanglement, as measured by concurrence, negativity, and quantum discord, indicates that the break up of an ordered phase by defects lies at the very heart of the process. A simple argument that directly links the scaling of concurrence to that of defect formation is contained in the analysis in Ref. [34]. Given the nonlocality of entanglement, one may have thought that the scaling of entanglement generation with $\tau$ should rather be governed by density correlators which, as shown in Refs. [39, 40], cannot be derived solely from the assumptions underlying KZM, and hence go beyond the scaling of defect density formation. Importantly, however, and as shown explicitly in Ref. [39] for the two-point density correlators, the relevant length and time scales that appear in these correlators do obey KZM scaling. This resolves a potentially thorny conceptual issue.
@@
The first steps in this direction were taken by Dutta et al. [50] who showed that an isolated system driven across a QPT by a noisy control field exhibits an “anti-Kibble-Zurek” (AKZ) behavior, whereby slower driving results in a higher density of defects. The authors argued that the KZM dynamics decouples from noise-induced effects, leading to an additive scaling formula for the density of defects in the limit of weak noise, $n\sim r\tau+c\tau^{-\beta}$, with $r$ the rate at which the presence of noise generates defects, and with $c$ a nonuniversal constant. When the first term dominates, the anti-Kibble-Zurek behavior becomes manifest, as fleshed out in later theoretical [51, 52, 53, 54, 55, 56] and experimental [57] studies (cf. Appendix A).
@@
We here address these questions in the setting of the TFI chain, the same model as used in the case studies in Refs. [50, 51]. Specifically, we analyze the two-spin entanglement generated by noisy quenches across the QPTs of the model, taking white noise as a representative, and added as a classical stochastic variable to the transverse field. Since one does not have access to individual noise realizations in an experiment, we consider the ensemble average over many realizations, with the average density matrix $\langle\rho\rangle$ emerging as the relevant quantity of interest. To obtain $\langle\rho\rangle$, we solve an exact master equation for the quench dynamics averaged over the noise distribution. By numerically solving the master equation, we then extract the average entanglement produced by an ensemble of noisy quenches. When two-spin entanglement is present, as measured by concurrence, we find that it scales logarithmically with the time scale $\tau$ when the quenches are slow, with an amplitude that depends on the intensity $\xi$ of the corresponding white-noise distribution. This is different from the behavior of the defect density in the presence of white noise as predicted by the anti-Kibble-Zurek additive formula in Refs. [50, 51]. As a consequence, the notion that entanglement and defect formation are intrinsically linked becomes less persuasive. A possible counterpoint is our finding that the time scale $\tau_{c}(\xi)$ beyond which no two-spin entanglement is produced scales as a power law of the noise intensity with the very same exponent as that of the optimal quench time for which the density of defects is minimized in a quantum annealing scheme [50, 51]. Whereas a direct, quantitative link between two-spin entanglement generation and defect formation seems unlikely in the presence of noise, this latter result hints at a possible connection, albeit more complex, between the two phenomena.
To prepare for our analysis, in the next section we set up the formalism and define the model. As a backdrop, we then revisit the simple case of a noiseless quench in Sec. III. This allows us to present a new result: Different from a noiseless symmetric “very-large-field” quench as discussed in Ref. [33] (i.e., with $h_{i}=-h_{f}\rightarrow$-$\infty$), we find that a generic quench generates two-spin entanglement also between nearest-neighbor spins. In Sec. IV, we turn to the main focus of our work, the case of noisy quenches. We take off from the exact master equation mentioned above, display the results from its numerical solution, and extract the scaling laws for the two-spin averaged entanglement as measured by concurrence - generated by slow noisy quenches across the quantum critical points of the model. Section V contains a brief summary and outlook. Mathematical details of our analysis can be found in the Appendix.
@@ II Model and formalism
[where $s^{\pm}_{n}=\sigma^{\pm}_{n}/2=(\sigma^{x}_{n}\pm i\sigma^{y}_{n})/2$], and introducing the discrete Fourier transforms $c_{k}=(1/\sqrt{N})\sum_{n}e^{ikn}c_{n}$ with $k=\pm(2j\!-\!1)\pi/N,j=1,2,\ldots N/2$, $H_{0}(t)$ can be expressed as a sum over decoupled mode Hamiltonians $H_{0,k}(t)$,
@@
set by the angle $\theta_{k}(t)=\arctan({\Delta_{k}}/{h_{0,k}(t)})/2$. In this basis, the Hamiltonian in Eq. (3) takes the diagonal form
@@
with $\varepsilon_{k}(t)=\sqrt{h^{2}_{k}(t)+\Delta_{k}^{2}}$. The eigenstates of $H_{0}(t)$ factorize into quasiparticle states labeled by $k$. Introducing $|0\rangle$ as the vacuum of the JW fermions $\{c_{k},c_{k}^{\dagger}\}$, these states can be written as
@@
In the case of a noiseless ramped quench we can write $h_{0}(t)\!=\!(t-t_{i})/\tau+h_{i}$, with $\tau=(t_{f}-t_{i})/(h_{f}-h_{i})$ the time scale of the quench, where $h_{i}\,(h_{f})$ is the initial (final) value of the ramp field at time $t_{i}\,(t_{f})$. To address the problem of noisy quenches, a direct approach would be to add a stochastic variable $\eta(t)$ to the ramp field, making the replacement $h_{0}(t)\rightarrow h(t)=h_{0}(t)+\eta(t)$ in the formulas above, compute the observable of interest - or, in our case, an entanglement measure - and then take an average over many noise realizations $\eta(t)$. In practice this is a cumbersome procedure, which, however, can be cut short by using an exact master equation, to be discussed in Sec. IV (see also Appendix C).
A standard choice for this class of applications is to take the noise to be Gaussian with vanishing mean, $\langle\eta({t})\rangle=0$, and with Ornstein-Uhlenbeck two-point correlations [61]:
@@
Here, $\tau_{n}$ is the noise correlation time and $\xi$ the noise amplitude for fixed $\tau_{n}$. For simplicity, and for comparison with the results in Refs. [50, 51, 52, 53, 54, 55, 62], we shall confine our attention to the common white-noise limit,
@@ III Entanglement generation from a noiseless ramped quench
In Fig. 1(a), we present our results for the concurrence $C_{\l ,\l +1}$ between two nearest-neighbor spins as a function of the endpoint quench field $h_{f}=h_{0}(t)$ at time $t$, with the quench starting at time $t_{i}$ with $h_{i}=-30$. Three graphs are plotted, for a slow ($\tau=500$), intermediate ($\tau=10$), and fast ($\tau=0.2$) quench, all for a chain with $N=200$ sites. The black dotted curve serves as a reference, showing $C_{\l ,\l +1}$ for the instantaneous ground state. As anticipated, and as shown in the inset of Fig. 1(a), this curve is in agreement with that known in the literature for $h_{0}(t)\geq 1/2$ [64, 65]. Its minimum at $h_{0}(t)=0$ reflects the multipartite entanglement of the zero-field Greenberger-Horne-Zeilinger (GHZ) ground state for a finite ferromagnetic TFI chain, implying a suppression of the two-spin entanglement.
@@
For intermediate values of the time scale, with $\tau$ in the interval $1\lesssim\tau\lesssim 100$, $C_{\l ,\l +1}$ exhibits oscillations, exemplified by the green curve in the inset of Fig. 1(a). The oscillations, originating from the phase factors of the off-diagonal elements $v_{0,k}(t)u_{0,k}^{\ast}(t)$, $v_{0,k}^{\ast}(t)u_{0,k}(t)$] of the density matrix $\rho_{0,k}(t)$ [14, 39], become more volatile in the lower range of the interval $1\lesssim\tau\lesssim 100$, but get dampened as $h_{0}(t)$ increases.
As seen from the blue graph for $\tau=0.2$, a further reduction of the time scale of the quench causes $C_{\l ,\l +1}$ to abruptly drop to zero after crossing the first quantum critical point at $h_{0}(t)=-1$. $C_{\l ,\l +1}$ then suddenly shoots up when crossing the second critical point at $h_{0}(t)=1$, with one more spike soon thereafter before vanishing identically. This striking behavior, common to all time scales $\tau\lesssim 0.5$, shows that a very fast quench across a QPT may completely wipe out the two-spin entanglement, instead triggering multipartite entanglement [cf. the instantaneous GHZ ground state at $h_{0}(t)=0$] or a mutually factorized two-spin state, with the latter possibly forming an entangled composite with other spins of the system.
@@
Figure 2: (Color online) Noise-averaged concurrence between two spins at time $t$ for a ramped quench of time scale $\tau=10$ with white noise of different intensities $\xi$ added to the transverse field of the TFI chain, with quench interval from $h_{i}=-30$ at $t=t_{i}$ to $h_{f}=h_{0}(t)$: (a) nearest-neighbor spins and (b) next-nearest-neighbor spins. System size: $N=200$.
@@
Figure 3: (Color online) Noise-averaged concurrence between two spins as a function of time scale $\tau$ for a ramped quench with added white noise of different intensities $\xi$, with quench interval from $h_{i}=-30$ to (a) $h_{f}=5$ for nearest-neighbor spins, and (b) $h_{f}=30$ for next-nearest-neighbor spins. System size: $N=200$.
Independent of the choice of time scale $\tau$, all quenches generate entanglement between neighboring spins that decays to zero when $h_{0}(t)\rightarrow\infty$, consistent with the analytical prediction by Sengupta and Sen [33]. These authors predicted a different type of behavior for the next-nearest-neighbor concurrence $C_{\l ,\l +2}$, with two distinct asymptotics for very large $h_{0}(t)$: $C_{\l ,\l +2}\neq 0$ $C_{\l ,\l +2}=0$] when $\tau>\tau_{0}\approx 1.96$ $\tau<\tau_{0}$]. Our results for $C_{\l ,\l +2}$, displayed in Fig. 1(b) as a function of the endpoint quench field $h_{0}(t)$ at time $t$, with the quench starting at time $t_{i}$ with $h_{i}=-30$, agrees with this prediction: By extrapolating to large $h_{0}(t)$, we find numerically that $1.9<\tau_{0}<2.0$.
While consistent with the prediction of the time scale $\tau_{0}$ [33], the overall behavior of $C_{\l ,\l +2}$ as shown in Fig. 1(b) is unexpected from the perspective of KZM. Only for slow quenches [cf. the red curve with $\tau=500$ in Fig 1(b)] is there a weak imprint on $C_{\l ,\l +2}$ that the quench has approached or crossed the critical point $h_{c}=-1$ (but not $h_{c}=1$). Besides this, values of $C_{\l ,\l +2}$ remain very small - not visible in Fig. 1(b) - for all $h_{0}(t)\lesssim 2$. This contrasts with the nearest-neighbor concurrence $C_{\l ,\l +1}$ which exhibits a dramatic increase when approaching the first critical point $h_{c}=-1$, supporting the conjecture that the entanglement generated reflects the KZM formation of defects when the dynamics turns nonadiabatic [19, 14]. One may naively have thought that the fast drop of $C_{\l ,\l +1}$ in the ferromagnetic phase $1<h_{0}(t)<1$ [cf. Fig. 1(a)] would influence $C_{\l ,\l +2}$ by a restructuring of the entanglement with closest further neighbor spins. But not so. Instead, the next-nearest-neighbor concurrence is driven by the postcritical dynamics when the magnitude of the transverse field becomes sufficiently large. It is interesting to note that for near-adiabatic evolution (very slow quench: $\tau=500$), $C_{\l ,\l +2}$ closely follows its instantaneous counterpart (static reference curve) only when $h_{0}(t)<h_{c}=-1$, as illustrated in Fig. 1(b). Contrary to expectations, its behavior is distinct from the reference curve for $h_{0}(t)>-1$. This finding, similar to that above for $C_{\l ,\l +1}$ with $h_{0}(t)\gtrsim 1$, reflects how entanglement generation in non-equilibrium systems can differ from its equilibrium counterpart, even during a nearly adiabatic evolution.